# Thermalization rates and quantum Ruelle-Pollicott resonances: insights from operator hydrodynamics

### v1

@@ Abstract
There are some outstanding questions and comments. First of all, we have made the strong assumption that operator growth in ergodic Floquet many-body systems is governed by the same biased diffusive hydrodynamics as in the RUC case. This is certainly expected McCulloch and von Keyserlingk (2022); Zhou and Nahum (2020) but has not been proved rigorously. One might expect that if this conjecture does hold, then the leading eigenvalue of $P_{\ell}\mathbb{U}_{F}$, and hence $\overline{g}$, should simply be determined by the butterfly velocity ($v_{B}$), and front diffusion constant ($D$) of the corresponding hydrodynamical process. This is actually false; we give a family of models of biased diffusion in which have identical $v_{B},D$, but which predict different values of $\overline{g}$sup . Thus $\overline{g}$ is, if it can entirely be determined by the operator hydrodynamical equations, necessarily sensitive to details of the hydrodynamics which would usually be considered irrelevant.
On the practical side, it would be useful to find an efficient numerical method for determining the effective stochastic matrix $T$ given an ergodic unitary evolution, perhaps using some of the ideas in Refs. McCulloch and von Keyserlingk, 2022; Zhou and Nahum, 2020. There are other questions related to fundamental questions in the physics of ergodicity in quantum systems. For example, can one show that if all local operators decay exponentially, then the unitary dynamics has to have a biased random walk description like it does in RUCs? How are the results modified for ergodic Hamiltonian evolution, where the only conserved quantity is energy?
Finally, it would be interesting to see if similar ideas can be applied to classical chaos. Specifically, is there a useful analogue of operator hydrodynamics in classical dynamics? Observables on the classical phase space are functions, which can be decomposed into Fourier components. Similar to how a single operator string evolves into a sum of operator strings with growing size, a classical observable may grow under ergodic dynamics into a complicated sum with increasingly high frequency Fourier components. It would be interesting to check, for simple models of classical chaos, if this kind of observable growth in Fourier space has a simple biased random walk description like operator size in RUCs, and if properties of this random walk can be used to compute RP resonances. Furthermore, it has been shown that in single-particle quantum maps, a coarse-grained quantum Frobenius-Perron operator has eigenvalues that converge to the corresponding classical RP resonances in the semiclassical limit followed by the limit of weak coarse-graining. In contrast, for quantum mechanical models whose classical limits are not chaotic, all the eigenvalues go back to the unit circle in the said limits. It will be interesting to see if such dichotomy exists for many-body quantum systemsNonnenmacher (2003); García-Mata and Saraceno (2004).
@@
Acknowledgements— We thank Sarang Gopalakrishnan, Tibor Rakovszky, Eric Schultz, Dries Sels, Sagar Vijay, and Zixia Wei for helpful conversations. C.Z. is supported by the Harvard Society of Fellows and the Simons Collaboration on Ultra Quantum Matter. CvK is supported by a UKRI Future Leaders Fellowship MR/T040947/2. This work was initiated and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 and a grant from the Alfred P. Sloan Foundation (G-2024-22395).
@@
Note added— We recently became aware of a parallel and independent work by Jacoby, Huse, and Gopalakrishnan Jacoby et al. which will appear on the same arXiv posting as this manuscript. Our conclusions appear to agree where they overlap.

### v2

@@ Abstract
Discussion.—There are some outstanding questions and comments. First of all, we have made the strong assumption that operator growth in ergodic Floquet many-body systems is governed by the same biased diffusive hydrodynamics as in the RUC case. This is certainly expected McCulloch and von Keyserlingk (2022); Zhou and Nahum (2020) but has not been proved rigorously. One might expect that if this conjecture does hold, then the leading eigenvalue of $P_{\ell}\mathbb{U}_{F}$, and hence $\overline{g}$, should simply be determined by the butterfly velocity ($v_{B}$), and front diffusion constant ($D$) of the corresponding hydrodynamical process. This is actually false; we give a family of models of biased diffusion in which have identical $v_{B},D$, but which predict different values of $\overline{g}$sup . Thus $\overline{g}$ is, if it can entirely be determined by the operator hydrodynamical equations, necessarily sensitive to details of the hydrodynamics which would usually be considered irrelevant.
On the practical side, it would be useful to find an efficient numerical method for determining the effective stochastic matrix $T$ given an ergodic unitary evolution, perhaps using some of the ideas in Refs. McCulloch and von Keyserlingk, 2022; Zhou and Nahum, 2020. It would also be interesting to understand how the results are modified in the presence of symmetries, where one can study the gap of the projected dissipative Floquet operator (projecting out the conserved quantities). Finally, it would be interesting to see if an analogue of operator hydrodynamics can be applied to classical chaos.
@@
Acknowledgements— We thank Sarang Gopalakrishnan, Tibor Rakovszky, Eric Schultz, Dries Sels, Sagar Vijay, and Zixia Wei for helpful conversations. C.Z. is supported by the Harvard Society of Fellows and the Simons Collaboration on Ultra Quantum Matter. CvK is supported by a UKRI FLF through MR/T040947/2 and MR/Z000297/1. This work was initiated and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 and a grant from the Alfred P. Sloan Foundation (G-2024-22395).
@@
Note added— We recently became aware of a parallel and independent work by Jacoby, Huse, and Gopalakrishnan Jacoby et al. (2024) which will appear on the same arXiv posting as this manuscript. Our conclusions appear to agree where they overlap.